OpenStudy (anonymous):

Evaluate the surface integral z dS where S is the part of the plane 4x+3y+z=12 that lies in the first octant.

OpenStudy (anonymous):

\[\int\limits_{?}^{?} z DS\]

OpenStudy (anonymous):

@Phi did one similar a while back, here's the link. http://assets.openstudy.com/updates/attachments/514bb348e4b05e69bface25f-phi-1364172797512-evaluatethesurfaceintegral.pdf

OpenStudy (anonymous):

@mathmale

OpenStudy (ipwnbunnies):

For a scalar function: \[\int\limits_{S}^{} f(x,y,z) dS = \int\limits_{R}^{} \int\limits_{}^{} f(x(u,v),y(u,v),z(u,v))*|r_{u} \times r_{v}| dA\]

OpenStudy (ipwnbunnies):

You have to rewrite the surface in parametric form. Sorry, if I'm no help. ;-;

OpenStudy (ipwnbunnies):

rewrite the equation of the surface*

OpenStudy (anonymous):

so z=12-4x-3y parametizes the surface...

OpenStudy (ipwnbunnies):

r(u,v) = <u, v, 12 - 4u - 3v>

OpenStudy (ipwnbunnies):

You can rewrite the function in terms of u and v now, or do that after the cross product.

OpenStudy (anonymous):

right or (x,y, 12-4x-3y)

OpenStudy (ipwnbunnies):

Uh, sure. Find the partial derivative of the function wrt u and v.

OpenStudy (ipwnbunnies):

Um, Idk what you did there.

OpenStudy (anonymous):

F wrt u=-4, wrt v=-3

OpenStudy (anonymous):

those are the partials^

OpenStudy (ipwnbunnies):

No the partial derivative of the vector function: r(x,y) = <x, y, 12 - 3x- 4y> With respect to x and y.

OpenStudy (ipwnbunnies):

We need to cross product rx and ry. Then find the magnitude.

OpenStudy (anonymous):

idk what you are saying?

OpenStudy (ipwnbunnies):

I know. Calculus sucks in typing.

OpenStudy (anonymous):

OH nvm. ok so <1,1, 12-3x-4y> right?

OpenStudy (ipwnbunnies):

No, we need partial derivative of x and y separately. rx = <1, 0, -3>

OpenStudy (mathmale):

Your (x,y, 12-4x-3y) appeals to me because it expresses three variables in terms of just two: x and y. Unfortunately, it's been some months since I last did problems of this nature, and I'd have to review that material to be of significant help. iPwnBunnies seems to be current and well informed, so I'd encourage you two to continue working together on this problem.

OpenStudy (ipwnbunnies):

Yes, it's a parametric form of the surface lol. Since for the surface, z is a function of x and y.

OpenStudy (ipwnbunnies):

Need it for the working definition of a surface integral.

OpenStudy (anonymous):

oh sorry I didn't understand what you were asking for

OpenStudy (anonymous):

ry= <0,1,-4>

OpenStudy (ipwnbunnies):

Oh crap, I think I mixed up the coefficients for the equation of the surface.

OpenStudy (anonymous):

yeah you did. but fixing it would be <1,0,-4> and <0, 1, -3>

OpenStudy (anonymous):

for rx and ry respectively

OpenStudy (ipwnbunnies):

z = 12 - 4x - 3y r(x, y) = <x, y, 12 - 4x - 3y> rx = <1,0,-4> ry = <0, 1, -3> Good. Can you cross product rx and ry, and find the magnitude?

OpenStudy (anonymous):

<4, -3, 1> magnitude of sqrt(16+9+1) or sqrt(26)

OpenStudy (ipwnbunnies):

Good, let's go back to the working definition. I'll just draw it.

OpenStudy (ipwnbunnies):

 We have the cross product done. Our function is not in terms of x and y yet doh. We can get it from the parametric form of the surface though.